OpenStudy (ccastudentgrade10):

FANANDMEDAL let f(x)=x-2 and g(x)=x^2-3x+2 find f(x)*g(x) and find f(x)-g(x) i have to do a bunch of these and i dont understand how. my teacher didn't give us an example problem and she's out of office today. i asked it before but no one finished helping me pls help!!

OpenStudy (mathmale):

Yes: You had this problem posted earlier, and I made a point of explaining what you had to do. What happened to those explainations? You could tag me if you want my attention.

OpenStudy (ccastudentgrade10):

i tagged you twice and you didn't respond?

OpenStudy (mathmale):

What questions did you have for me then?

OpenStudy (ccastudentgrade10):

i tagged you again on the other

OpenStudy (mathmale):

Given f(x) and g(x), how do you express their product? Given f and g, how do you express their difference?

OpenStudy (mathmale):

What did you want from me at that point?

OpenStudy (mathmale):

Is your previous post still o0pen?

OpenStudy (ccastudentgrade10):

oh its closed but what i asked was 1. so is it x^2 - 2 and 2. i just dont know where to start with the multiplication

OpenStudy (mathmale):

(a+b)(c+d)=ac+ad+ ? + ? This is an example of the multiplication of binomals.

OpenStudy (mathmale):

I applied the Distributive Property of Multiplication, multiplying (c + d) by a first: ac + ad. Now multiply (c + d) by b:

OpenStudy (mathmale):

I'll work with you on this if you'd please give this your full attention.

OpenStudy (ccastudentgrade10):

im sorry my computer crashed @mathmale

OpenStudy (ccastudentgrade10):

someone help please

zepdrix (zepdrix):

\[\large\rm f(x)=(x-2)\qquad\qquad\qquad g(x)=(x^2-3x+2)\]So then,\[\large\rm f(x)\cdot g(x)=(x-2)(x^2-3x+2)\] Mmm k this is going to be a bit of a doozy, let's see...

zepdrix (zepdrix):

We would start by giving this x in the first set of brackets, to each term in the next set of brackets.